A helicopter is descending vertically at a constant speed of 3 ms-1. A sandbag is released from the helicopter. The sandbag hits the ground 5.0 s later. What was the height of the helicopter above the ground at the time the sandbag was released?


1
Expert's answer
2021-03-08T07:27:42-0500

Let's take the downwards as the positive direction. Then, we can find the height of the helicopter above the ground at the time the sandbag was released from the kinematic equation:


"y=v_0t+\dfrac{1}{2}gt^2,""y=3\ \dfrac{m}{s}\cdot5\ s+\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(5\ s)^2=137.5\ m."